# Feature Calibration enhanced Parameter Synthesis for CLIP-based Class-incremental Learning

### v2

@@ IV-B1 Results on ImageNet100
Table I shows the results of our method on ImageNet100 under five different incremental settings. It is clear that the non-traditional CIL method significantly outperforms traditional CIL methods, achieving a substantial performance gap. It indicates that CLIP holds great potential in the field of CIL, likely due to the positive impact of incorporating text features. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 2.3%, 3.21%, 3.59%, 4.15%, and 4.77% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. It suggests that our method is not solely reliant on the inherent generalization ability of CLIP but also the effectiveness of our proposed FC and PS techniques. Compared to DualPrompt [15] and CODA [16], our method performs better, demonstrating that optimizing and improving the text prompt space alone is insufficient. It is also necessary to enhance the use of visual features. Under the B0 experimental setting, our method achieved an average accuracy of 88.04%, 87.62%, and 88.19%, which are at least 0.45%, 0.68%, and 0.9% higher than those of other methods, showcasing the superiority of our approach. From the last metric, our method consistently ranks at the top, maintaining around 80% with slight fluctuations, whereas other method’s last values are much lower. It indicates that our method exhibits strong resistance to forgetting, confirming the rationality of incorporating parameter decomposition, fusion, and dynamic distillation ideas in our approach.
@@ IV-B2 Results on CIFAR100
Table II shows the results of our method on CIFAR-100 under five different incremental settings. It is evident that the non-traditional CIL method outperforms traditional methods, with an even larger and more pronounced performance gap compared to its performance on the dataset ImageNet100. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 10.77%, 11.06%, 11.44%, 12.25%, and 12.49% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. These gains are nearly 2 to 3 times greater than those observed on the dataset ImageNet100, further validating the powerful capabilities of FC and PS. Under the B0 experimental setting, our method achieves an average accuracy of 86.70%, 86.06%, and 85.45% across different incremental task settings, which are at least 1.58%, 1.18%, and 0.26% higher than those of other methods. It demonstrates the excellence of our approach. The metric last of our method is consistently above 79.50% in various experimental settings, a level that is nearly unattainable for other methods. These highlight the significant advantage of our method in addressing the forgetting problem.
@@ IV-C1 Module Ablation
ID FC PS Distill Avg Last
1 84.98 75.40
2 85.82 73.78
3 88.16 80.02
4 88.19 80.32
Table III presents the results of ablation experiments on different modules of our proposed method. The FC module contributes a 0.84% improvement in accuracy, indicating that retaining the original CLIP features is crucial. It prevents the potential distortion of CLIP original features caused by direct fine-tuning. However, the last metric decreases by 1.62%. It suggests that FC somewhat exacerbates forgetting to a certain extent, while plays a significant role in learning new knowledge. The PS module, which uses QR decomposition to retain old knowledge, improves accuracy by 2.34% and the last metric by 6.24%. It demonstrates that the module not only positively impacts the learning of new knowledge but is also highly effective in addressing the forgetting problem. It indicates that decomposing and fusing the adapter’s parameter matrix is an efficient approach, balancing the learning of new knowledge while preserving old knowledge. The dynamic distillation module plays a more prominent role in mitigating the forgetting problem.
@@ IV-C2 Effectiveness of Loss Function
In Eq. 9 , we primarily use cross-entropy loss and dynamic distillation loss, with the parameter $\lambda$ incorporated into the loss calculation. As shown in Table III, when we only use cross-entropy loss, the last accuracy of the method is 80.02%. When we introduce $\lambda$ and distillation loss, the last accuracy improves by 80.32%. Additionally, the $\lambda$ we introduce represents the ratio of old classes to all classes and is dynamically adjusted. When $\lambda$ is a fixed value, the accuracy decreases to some extent. It indicates that the dynamic design of $\lambda$, which considers the proportion of old and new classes, allows the loss function to adapt dynamically, leading to improved performance.
@@ IV-C3 Sensitivity Analysis on $\alpha$ and $\beta$
When preserving the original CLIP features, $\alpha$ represents the degree of preservation of the original features. It requires a trade-off between the CLIP features and the enhanced features. As shown in Fig. 4(a), the size of $\alpha$ has a certain impact on the accuracy, indicating that whether and to what extent the original CLIP features are preserved is crucial. Compared to $\alpha$, the size of $\beta$ has a more direct and profound impact on the accuracy of the model. An inappropriate $\beta$ can have a detrimental effect on the performance of model in Fig. 4(b), further emphasizing the importance of the decomposition and fusion of the adapter parameters.
@@ V Conclusion
This paper proposes a novel CIL method called FCPS which integrates visual and textual features into the continual learning framework. A specific parameter update mechanism is designed to address the issue of feature distortion caused by directly fine-tuning vision-language models, effectively maintaining the stability and generalization of the model. Additionally, parameter update and integration approach is introduced for different task stages to mitigate catastrophic forgetting. Experimental results demonstrate the superiority of this method. Some parameter values were manually determined, and future work will focus on developing dynamic adaptive update methods while also exploring few-shot CIL to address the limited availability of new class data in real-world scenarios.

### v4

@@ IV-B1 Results on ImageNet100
Table I presents the performance of our proposed method on the ImageNet100 dataset under five distinct incremental learning settings. It is evident from the results that our approach, which deviates from conventional CIL paradigms, significantly outperforms traditional methods. The observed performance gap is substantial and consistent across all settings, highlighting the effectiveness of integrating CLIP into the CIL framework. This performance boost is largely attributed to the incorporation of text features, which enhance the semantic alignment between visual and textual modalities. These results underscore the promising potential of CLIP in incremental learning scenarios, where maintaining performance across sequential tasks is notoriously challenging. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 2.3%, 3.21%, 3.59%, 4.15%, and 4.77% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. It suggests that our framework does not rely solely on the strong generalization capabilities of pre-trained CLIP representations. Instead, the gains reflect the effectiveness of our novel FC and PS strategies, which are specifically designed to address the challenges of catastrophic forgetting and representation shift in CIL. Compared to DualPrompt [22] and CODA [23], our method consistently achieves superior results. While DualPrompt and CODA focus primarily on optimizing the textual prompt space, our findings suggest that improvements in prompt engineering alone are insufficient for robust CIL performance. Our approach leverages both textual and visual modalities by enhancing visual feature utilization, demonstrating that a more holistic integration of multimodal representations is critical. Under the B0 experimental setting, our method achieve average accuracy of 88.04%, 87.62%, exceeding the next-best methods by margins of at least 0.45%, 0.68%, and 0.9%, respectively. These consistent gains further validate the robustness and superiority of our proposed techniques. Additionally, when evaluating the performance based on the final task accuracy—a critical indicator of a model’s resistance to forgetting—our method maintains a performance level around 80%. In contrast, other methods exhibit significantly lower final accuracies, indicating greater susceptibility to catastrophic forgetting. These findings confirm that our method offers a well-balanced and effective solution for CIL. The integration of parameter decomposition, feature fusion, and dynamic distillation not only mitigates the drawbacks of previous approaches but also introduces a resilient framework that adapts well to evolving class distributions.
@@ IV-B2 Results on CIFAR100
Table II presents the performance evaluation of our proposed method on the CIFAR-100 dataset under five distinct incremental learning scenarios. As shown, non-traditional CIL approach significantly outperforms existing traditional methods, exhibiting an even more pronounced performance gap compared to the results observed on the ImageNet100 dataset. This substantial improvement underscores the generalizability and robustness of our approach across different datasets, including those with smaller resolution and less semantic complexity like CIFAR-100. Compared to the zero-shot performance of CLIP, our method achieves average incremental accuracy gains of 10.77%, 11.06%, 11.44%, 12.25%, and 12.49% on the B0Inc5, Inc10, Inc20, and B50Inc5, Inc10 incremental task settings, respectively. These gains are approximately two to three times greater than those observed on ImageNet100, further confirming the effectiveness of the proposed FC and PS mechanisms. Under the B0 experimental setting, our method achieve average accuracy of 86.70%, 86.06%, and 85.45% across different incremental task settings. These results are consistently higher than those achieved by competing methods, providing strong evidence for the robustness and superiority of our approach. Furthermore, the last metric, which indicates the final task performance and serves as a key measure of a model’s resilience to catastrophic forgetting, remains consistently above 79.30% across all experimental settings. In contrast, other methods are nearly unattainable to reach this level, often experiencing significant performance degradation over time. This consistent and high final performance highlights the effectiveness of our method in mitigating the forgetting problem—a central challenge in CIL. Overall, the outstanding performance of our method on CIFAR-100, demonstrates that our approach not only scales well across datasets but also provides a powerful and general framework for CIL.
@@ IV-C1 Module Ablation
FC PS Distill $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$ $A_{10}$ (Last) Avg
1 97.20 95.20 91.93 88.75 83.56 81.80 81.00 78.89 76.10 75.40 84.98
2 96.6 97.00 92.43 90.80 83.64 84.13 82.27 80.97 76.56 73.78 85.82
3 96.80 96.60 93.87 91.70 87.20 85.52 85.03 84.60 80.27 80.02 88.16
4 96.80 96.10 94.27 91.50 87.16 85.43 85.09 84.82 80.42 80.32 88.19
Table III presents the results of our ablation study, which evaluates the individual contributions of each module in our proposed method. Overall, our method demonstrates stable performance throughout all incremental stages. Meanwhile, the table shows the role of FC, PS, and distillation in enhancing model performance and mitigating catastrophic forgetting. The inclusion of the FC module leads to a 0.84% increase in average incremental accuracy, demonstrating that preserving the integrity of the original CLIP features is highly beneficial. FC prevents the distortion of pre-trained visual embeddings that often results from direct fine-tuning, thereby retaining the strong generalization capacity of CLIP. However, we observe a 1.62% drop in the last metric when FC is used alone, indicating that while it facilitates the acquisition of new knowledge, it may introduce a trade-off in terms of long-term memory retention. This suggests that FC enhances short-term task performance but slightly exacerbates forgetting over time if used in isolation. PS shows a more substantial contribution. It uses QR decomposition to retain old knowledge, improves accuracy by 2.34% and the last metric by 6.24%. These results clearly indicate that PS not only supports effective learning of new tasks but also plays a critical role in preserving previously learned knowledge. This module leverages the orthogonality principle in QR decomposition to facilitate better knowledge retention across tasks. The improvement in the last metric suggests that decomposing and fusing the adapter’s parameter matrix is a highly effective strategy, achieving a delicate balance between plasticity (learning new tasks) and stability (retaining old knowledge). The dynamic distillation module plays a more prominent role in mitigating the forgetting problem. By adaptively distilling knowledge from earlier task models and dynamically updating the student model, this module reduces the drift of representations and promotes consistency across learning stages. Its contribution becomes especially evident in the sustained high values of the last metric when combined with other modules, underscoring its vital role in preventing catastrophic forgetting. Overall, the ablation study validates the effectiveness of each individual module in our framework. While FC aids in preserving feature quality and learning new information, PS ensures that old knowledge is not overwritten, and dynamic distillation offers a strong regularization mechanism for incremental learning. The synergy among these components forms the foundation of our robust and adaptive CIL framework.
@@ IV-C2 Effectiveness of Loss Function
In Eq.9, our loss function is composed of two key components: the standard cross-entropy loss, which facilitates the learning of new class information, and a dynamic distillation loss, which serves to preserve knowledge from previously learned tasks. These two components are weighted by a scalar parameter $\lambda$, which plays a crucial role in balancing the influence of old and new class information during training. As evidenced by the ablation results in TableIII, this design significantly impacts the model’s ability to mitigate forgetting and retain prior knowledge. When only the cross-entropy loss is used—without incorporating any distillation mechanism—the final task accuracy (last) is 80.02%. This baseline reflects the typical performance of a model focused solely on learning new information without explicit mechanisms to retain prior knowledge. However, when we introduce the distillation loss alongside the cross-entropy loss, and dynamically modulate their relative contributions using the parameter $\lambda$, the last metric improves to 80.32%. Although the numerical gain may appear modest, it demonstrates that even a small degree of incremental knowledge transfer can contribute meaningfully to long-term retention in CIL. The parameter $\lambda$ is not static, rather, it is dynamically computed based on the ratio of old classes to the total number of classes seen so far in the incremental learning process. This adaptive strategy ensures that as the task progresses and the balance between old and new classes changes, the loss function remains sensitive to the shifting importance of knowledge retention versus acquisition. In contrast, when $\lambda$ is fixed across all learning stages, the model’s accuracy decreases. This drop illustrates the limitation of static weighting schemes in dynamic learning environments. The results clearly indicate that the dynamic design of $\lambda$ contributes to better performance by providing an adaptive mechanism that reflects the class distribution at each stage. As the number of classes increases, and the potential for forgetting previously learned classes grows, the dynamic $\lambda$ ensures that the distillation loss is appropriately weighted to preserve old knowledge. Conversely, in early stages with fewer old classes, the model is allowed greater flexibility to focus on new knowledge acquisition. This adaptability allows our loss formulation to strike an effective balance between stability and plasticity, which is essential in incremental learning settings. Overall, the incorporation of a dynamically adjusted $\lambda$ improves the effectiveness of the distillation process. This design choice further reinforces the robustness and adaptability of our proposed framework in addressing the challenges of catastrophic forgetting.
@@ IV-C3 Sensitivity Analysis on $\alpha$ and $\beta$
In our framework, preserving the original CLIP features is a key design consideration, and we introduce the parameter $\alpha$ to control the degree of feature preservation. Specifically, $\alpha$ serves as a balancing coefficient that governs the trade-off between the original frozen CLIP features and the newly enhanced features derived from task-specific adaptations. As illustrated in Fig. 6(a), variations in the value of $\alpha$ yield noticeable changes in model accuracy, highlighting the sensitivity of the model’s performance to this balance. This observation underscores the importance of carefully calibrating how much of the pre-trained knowledge is retained versus how much is adapted to new tasks. When $\alpha$ is too small, the model may underutilize the rich semantic representations encoded in CLIP, leading to suboptimal generalization. Conversely, when $\alpha$ is too large, the model may become overly reliant on static features, limiting its ability to adapt to new class distributions in the incremental learning setting. Thus, selecting an appropriate value for $\alpha$ is essential to achieving a balanced integration of pre-trained and task-specific knowledge. In contrast to $\alpha$, the parameter $\beta$, which controls the influence of the decomposed and fused adapter parameters, has an even more direct and pronounced effect on model performance. As depicted in Fig. 6(b), improper tuning of $\beta$ can significantly degrade accuracy, revealing the sensitivity of the model to the fusion mechanism applied to the adapter’s parameters. This further emphasizes the critical role of parameter decomposition and fusion in our approach. The adapter module, which operates as a lightweight, trainable bridge between frozen CLIP representations and the downstream incremental tasks, relies heavily on the proper scaling and integration of its components. $\beta$ modulates the contribution of these components, and an ill-suited value can disrupt the representational harmony between old and new knowledge, ultimately hindering the model’s capacity to learn incrementally without forgetting. The experiments shown in Fig. 6 collectively reveal that both $\alpha$ and $\beta$ are vital hyperparameters, but they influence different aspects of the learning process. While $\alpha$ governs the trade-off between stability (retaining CLIP’s pre-trained knowledge) and plasticity (learning task-specific enhancements), $\beta$ directly affects the internal representational learning within the adapter. Optimal performance is achieved only when both parameters are carefully tuned, enabling a synergistic interaction between feature preservation and adaptable parameterization.
@@ V Conclusion
This paper proposes a novel CIL framework called FCPS, which integrates visual and textual features within a incremental learning paradigm. To address the feature distortion caused by directly fine-tuning VLMs such as CLIP, FCPS introduces FC to preserve the original semantic structure of pre-trained features. This design helps maintain the model’s generalization ability and stability across tasks. In addition, FCPS employs a stage-specific parameter integration strategy to mitigate catastrophic forgetting. By adaptively decomposing and fusing parameters at different task stages, the model can effectively balance the learning of new knowledge with the retention of previous knowledge. This approach enables FCPS to achieve strong performance in both incremental accuracy and long-term knowledge retention. Experimental results on benchmarks like ImageNet100 and CIFAR-100 demonstrate that FCPS outperforms traditional and recent CIL methods, confirming the advantages of multi-modal feature integration and the proposed training strategies. While some hyperparameters in FCPS were manually set, future work will focus on developing adaptive parameter tuning methods. Additionally, we aim to extend FCPS to few-shot CIL scenarios, which better reflect real-world conditions where new class data is limited.